A 1.10 inch diameter pipe is required to carry 47.5 gpm. What flow velocity is required in ft/s?
Answer #1

pipe = 7(1-10)² incha Cross-section area of A = .346 inch 475 gram per minute= = .792 g/sec 475 g/ = -792,2025 = .792. com y= (-3942.79 7 cmifleec - 0.048 an incro/sec u= .048 inch / sec = .138 inch/sec =(1138/12) ft/sec = .0115ft/see €11.5X15 #sec